OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 2, y = t2 - 4

OpenStudy (anonymous):

i legit just had this

OpenStudy (anonymous):

isolate the t in the first equation.

OpenStudy (anonymous):

so t=sq rt x?

OpenStudy (anonymous):

t = sqrt x/2?

OpenStudy (psymon):

\[t = \sqrt{x-2}\]

OpenStudy (anonymous):

t=sqrt x -2

OpenStudy (anonymous):

now plug in the equation you just go into the "t"" in the "y=" equation

OpenStudy (anonymous):

so then y = sqrt x-2- 4

OpenStudy (anonymous):

the second equation is supposed to look like, right? \[y=t^2-4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y = (sqrt x-2)^2 - 4

OpenStudy (anonymous):

yes, but since the square cacels out the square root, you will end up with y=x-2-4

OpenStudy (anonymous):

oh, so y = x-6

OpenStudy (anonymous):

yup, so did that help?

OpenStudy (anonymous):

yes! thank you

OpenStudy (anonymous):

no problem!

OpenStudy (tkhunny):

x = t2 + 2, y = t2 - 4 Really? Square roots and substitution? \(t^{2} = x-2\) \(t^{2} = y+4\) \(x-2 = y+4\)